# Black Holes as Condensation Points of Fuzzy Dark Matter Cores

### v1

@@ Abstract
We simulate the formation of Fuzzy Dark Matter (FDM) cores in the presence of a Black Hole (BH) to explore whether BHs can serve as seeds for FDM core condensation. Our analysis is based on the core-condensation via the kinetic relaxation process for random initial conditions of the FDM. We show that in general the BH merges with pre-collapsed mini-clusters and once they share location the BH oscillates within the core. The condensation takes place around the black hole and the FDM acquires a density profile consistent with the density of the stationary solution of the FDM+BH eigenvalue problem in average. The central density of the resulting core depends on the mass of the BH, which due to its motion relative to the FDM cloud produces a smaller time-averaged densities for bigger BH masses, which lead to a new diversity of central FDM core densities. Our results indicate that BHs can indeed act as focal points for FDM core condensation. As a collateral result, for our analysis we revised the construction of stationary solutions of FDM+BH and found a phenomenological formula for the FDM density that can be used to fit FDM cores around BHs.
@@
Fuzzy Dark Matter (FDM) assumes dark matter is an ultralight boson with mass of order $10^{-23}-10^{-21}$eV, whose behavior differs from that of cold dark matter at galactic scales, in particular it promotes core formation and shows wave-like phenomena that produce a particular distribution of matter at galactic core and halo scales Chavanis (2015); Niemeyer (2020); Hui (2021); Ferreira (2020). The structures formed by this particle are cores surrounded by envelopes with profiles similar to those of the CDM, being the core an essential fingerprint of the model determined from structure formation simulations (e.g. Schive et al. (2014a); Mocz et al. (2017); Veltmaat et al. (2018); Mocz et al. (2019); May and Springel (2021); Schwabe and Niemeyer (2022)).
@@
Core formation has been studied in more detail at small scale simulations in basically two scenarios, one is the multi-merger of cores, for example in Schive et al. (2014); Mocz et al. (2017); Schwabe et al. (2016); Álvarez-Rios et al. (2023); Álvarez-Rios and Guzmán (2024) that leads to a simple construction of core-halo structures. Another approach uses the kinetic relaxation as explanation and simulation of core formation out of random initial conditions Levkov et al. (2018); Chen et al. (2021). These two methods help studying the core condensation process of bosonic configurations, their mass-growth and the envelope profile of their halos using small box simulations Eggemeier and Niemeyer (2019); Chen et al. (2023); Purohit et al. (2023); Chen and Zhang (2024).
@@
An essential ingredient introduced in FDM phenomenology is the presence of black holes and their behavior within FDM cores. For example, in Chavanis (2019, 2020) the FDM core density profile properties are studied under various regimes of the boson gas and scenarios that include a black hole, in Wang and Easther (2022) the interaction of FDM and the BH is studied, in particular the dynamical friction and the drag of the wake behind the hole while moving, in Boey et al. (2024) also the dynamical friction of FDM cores in BH is studied in merger scenarios, in Koo et al. (2024) also the binary black hole merger moving inside FDM cores is studied in the context of the final parsec problem of the merger. In gravitational wave related contexts, the binary black hole within FDM is studied to estimate the extraction of angular momentum due to dynamical effects of clumps of dark matter within the core Bromley et al. (2024). In El-Zant et al. (2020) the authors study the ejection of SMBHs due to the superposition of modes and the accumulation of random walk effects within FDM halos. Also in Lancaster et al. (2020) the motion of a massive point particle in a FDM environment that includes its granules is studied, while in Boey et al. (2024) the collision between an FDM core and a SMBH is analyzed. These studies consider the black hole to be Newtonian, and focus mainly in dynamic effects of moving black holes. In a relativistic context, the coexistence and phenomenology of black hole with ultralight bosonic dark matter has also been analyzed, for example in Cardoso et al. (2022a); Traykova et al. (2023); Ravanal et al. (2023); Boudon et al. (2023); Traykova et al. (2023) various effects of black hole dynamics on the scalar field are studied, including dynamical friction, while in Bamber et al. (2023); Aurrekoetxea et al. (2024a) the analysis centers on the potential detection of ultralight dark matter via effects on gravitational wave phenomenology, including the case with self-interaction Aurrekoetxea et al. (2024b). In cosmological scenarios, studies include for example the nucleation of scalar clouds Hertzberg et al. (2020).
@@
In this work we study the effects of a BH during the FDM core-condensation, for which we follow the kinetic relaxation in Levkov et al. (2018); Chen et al. (2021), that uses random initial conditions, so that the granularity of the FDM distribution develops in the presence of the BH since initial time. In our analysis we use a Newtonian description of the BH and add its gravitational potential to the Scrödinger-Poisson system of equations that rules the dynamics of the FDM+BH system, while we ignore the partial accretion of wave dark matter, which depends on the wave-length of wave-packets approaching the black hole, as demonstrated with full non-linear numerical relativity in Guzmán and Lora-Clavijo (2012); Guzmán et al. (2021), as well as other non-linear accretion effects in non-symmetric scenarios Cardoso et al. (2022b).
@@
The system of equations that rule the dynamics of the FDM+BH system is the following Schrödinger-Poisson system of equations:
@@
where $m$ is the boson mass, $\rho=m|\Psi|^{2}$ is the bosonic gas density, $\bar{\rho}$ its mean density, $V$ is the gravitational potential due to the FDM and the BH, while $V_{FDM}$ is the potential only due to the FDM; we approximate the BH with a density distribution $\rho_{BH}=M_{MH}\delta(\vec{x}-\vec{x}_{BH})\simeq CM_{MH}e^{-|\vec{x}-\vec{x}_{BH}|^{2}/2\epsilon^{2}}$, which is an approximate Dirac delta distribution, where $C$ is such that the integral of the distribution is the black hole mass $M_{BH}$; the total density of the system is $\rho_{T}=\rho+\rho_{BH}$ and finally, $\vec{x}_{BH}$ is the position of the black hole, that reacts to the gravity of the FDM.
@@
In order to solve the SP system above, we use the transformations $t=t_{0}\tilde{t},\vec{x}=x_{0}\tilde{\vec{x}},V=V_{0}\tilde{V},\Psi=\Psi_{0}\tilde{\Psi}$ and $\rho=\rho_{0}\tilde{\rho}$, where $t_{0}=\frac{x_{0}^{2}m}{\hbar}$, $V_{0}=\left(\frac{\hbar}{mx_{0}}\right)$, $\Psi_{0}=\frac{\hbar}{\sqrt{4\pi Gm^{3}}x_{0}^{2}}$, $\rho_{0}=\frac{\hbar^{2}}{4\pi Gm^{2}x_{0}^{4}}$, that leave the SP system in dimensionless code units that only depend on the length scale parameter $x_{0}$. We solve the dimensionless problem using the code CAFE-FDM Álvarez-Rios and Guzmán (2022), that implements a pseudo-spectral method to discretize spatial derivatives, an RK4 scheme for the evolution of the wave function, and the two Poisson equations are solved using the FFT method. All simulations are carried out in a periodic cubic domain of side $L$ in code units, with resolution $\Delta x$. For the BH distribution we use $\epsilon=0.1\Delta x$ because it gives similar results as in Boey et al. (2024) for non-periodic domains, where a potential for the BH $V_{\bullet}=-\frac{GM_{BH}}{\sqrt{max(r^{2},\epsilon^{2})}}$ is used within Schrödinger equation, but cannot be straightforwardly implemented in a periodic domain. As a test of our implementation, we evolve the eigen-solutions of the stationary FDM+BH configurations developed in Davies and Mocz (2020) and verify they are stable.
@@
Initial conditions. The kinetic relaxation leading to core condensation uses random initial conditions that eventually lead to the collapse of overdensities that in turn promote the bosonic gas condensation. The systematic study of this collapse process is understood from local simulations defined in Levkov et al. (2018); Chen et al. (2021), where various distributions in the momentum domain are proposed, that give similar results. In our analysis we use a Gaussian distribution $\Psi(\vec{p})=Ae^{-0.5p^{2}}e^{iS}$ in the momentum space, with $S$ a random phase in the range $[0,2\pi]$ at each point of the momentum space, where $A$ is a normalization factor. In order to validate our methods and code, we run one of the simulation in Chen et al. (2021), specifically a box of side $L=18$, with resolution $\Delta x=L/128$ and total FDM mass $M_{FDM}=M=1005.3$ and the results are shown in Appendix A as a test. We use these settings as default in our simulations involving BHs.
@@
Simulations. We performed 32 simulations with the same type but different initial conditions for the FDM and various black hole masses $M_{BH}=M/256,M/128,M/64,M/32$, which lie within the range of masses used in studies of dynamical friction Wang and Easther (2022) and those of PBHs used for nucleation of axion stars in Hertzberg et al. (2020). Different initial seeds for initial conditions lead to different evolutions of the FDM, however we illustrate the generic evolution of the system using the same particular seed to generate the random phase of FDM in Fourier space that in turn generates the initial conditions of the FDM, in the simulations with the four black hole masses. The evolution of the four cases is presented in Figure 1, that contains snapshots of the FDM density and the Black Hole position. At early times by $t\sim 7$ a minicluster forms near the bottom-right corner, later on this cluster merges with the black hole, following different trajectories for each black hole mass, and finally the condensation takes place at the black hole location. Animations of these simulations appear as Supplementary Material Palomares-Cávez et al. .
@@
Figure 1: Snapshots of the density projected on a plane that contains the position of the black hole and is parallel to the $xy-$plane of the numerical domain. The black hole position is represented with a yellow cross at different times for simulations with $M_{BH}=M/256,M/128,M/64$ and $M/32$. The evolution illustrates the formation of a minicluster near the bottom right corner. Later on the minicluster and the BH merge and start traveling together while the BH oscillates with respect to the core that can be better seen in the movies in the supplementary material. By $t\sim 60$ the FDM approaches condensation as seen in the analysis below.
@@
Condensation. In order to see if condensation happens, we monitor the evolution of $\rho_{max}$ and the results are shown in Figure 2, where we notice that the condensation takes place, the density grows with a power law of $t$ starting after $t\sim 30$, which indicates the beginning of condensation, however later on the maximum density decreases or stabilizes instead of growing up as in the pure FDM scenario in Levkov et al. (2018); Chen et al. (2021), that we use as test in Appendix A. The decrease of $\rho_{max}$ is smaller for smaller black hole masses and can be understood in terms of the relative motion between the BH and the core, namely, the more massive the black hole is, the more FDM mass drags with it during relaxation. The black hole motion drags the FDM around and does not allow a quiet condensation, instead it is constantly revolving the FDM preventing the central density from growing, from becoming more compact. In this sense the Black Hole can act as an entity that decreases the final central FDM core density.
@@
Figure 2: Evolution of $\rho_{max}$ for the various black hole masses $M_{BH}=M/256,M/128,M/64,M/32$. Notice that for the less massive BHs the density continues growing until $t\sim 60$ where it stabilizes. A difference between small BH masses and the case $M_{BH}=0$ shown in Fig 7, is that $\rho_{max}$ stabilizes with a BH, whereas it keeps growing when there is not BH in this time window, although as we will see below, the central density in average for the cases $M_{BH}=0,M/256$ and $M/128$ are very similar. For the most massive BHs the maximun density decreases, which is an indication of the effects of the presence of the BH. For comparison of the evolution without BH see the simulations in Appendix A.
@@
Relative motion. In Figure 3 we show the distance from the point of maximum density to the Black Hole, for the four values of $M_{BH}$. The non-linear behavior and the dynamics of the granularity lead to behavior that does not have a trend, since the motion of the black hole can be reheated differently since the different four mergers occur under different conditions, at different places and with different merging velocities, as can be seen in the animations.
@@
Another part of the analysis is the velocity of the BH. In Figure 4 we show the $x-$component of the black hole’s velocity, which shows the oscillatory behavior studied under smoother FDM distributions in Boey et al. (2024). Moreover, for the cases $M_{BH}=M/128$ and $M/64$ the reheating pulses can be observed in the velocity as also described in Boey et al. (2024). Unlike in Boey et al. (2024) the oscillations are less regular because we do not assume an initially smooth spherically symmetric core, but deal with the core resulting from the collapse of FDM, with all its granular multipolar components and time dependence, since the formation time.
@@
Figure 4: The $x-$component of the BH velocity, for $M_{BH}=M/256/,M/128,M/64,M/32$. These oscillations are consistent with those proposed in Boey et al. (2024) for mergers of smooth cores and BHs. In our case the core has all its granules and time dependency since the beginning. Notice that for $M_{BH}=M/128,M/64$ the modulation of the velocity amplitude is consistent with the reheating of the BH motion. The behavior of the $y$ and $z-$components of the velocity is similar. The animations in the Supplementary Material illustrate the whole picture of the process.
@@
Core fitting. Even though there is dynamics, we fit the resulting density profile of the FDM, not only at a fixed time but averaged over a time window using a phenomenological model for the density, that mimics the solution of the FDM-BH stationary eigen-problem in Davies and Mocz (2020), as described in Appendix B; the FDM density fitting uses formulas (7)-(9). The result shows the possibility to fit the configuration with an equilibrium profile of the FDM-BH eigen-problem. The fittings are presented in Figure 5, that we obtain from averaging the density profile in the time window $t\in[70,100]$, which is a lapse when, according to Fig. 2, the maximum density profile ends growing during condensation at least for the smallest black holes. Notice that the central core density is smaller for bigger BH mass, whereas the central core density converges to the case without BH when $M_{BH}$ becomes smaller, in fact the cases $M_{BH}=M/256$ and $M/128$ have a profile similar already to that of the case $M_{BH}=0$.
@@
Figure 5: Density fitting of the angularly and time averaged density of FDM around the BH, using the phenomenological formula (7), for the simulations with $M_{BH}=M/256,M/128,M/64,M/32$. The time interval used to average density profiles is $t\in[70,100]$. Notice that the more massive the black holes, the smaller the central core density, which is consistent with the notion of a Black Hole moving within the core while dragging FDM around and decreasing density, which is consistent with the results in Figure LABEL:fig:rhomaxMBH. We enclose the case for $M_{BH}=0$ to show that when the mass of the BH approaches zero, the core approaches the case without BH.

### v3

@@ Abstract
We simulate the formation of Fuzzy Dark Matter (FDM) cores in the presence of a Black Hole (BH) to explore whether BHs can serve as seeds for FDM core condensation. Our analysis is based on the core-condensation via the kinetic relaxation process for random initial conditions of the FDM. In a generic scenario the BH merges with a pre-collapsed mini-cluster formed in a random location, once they share location the core-condensation starts withe the FDM density centered at the black hole that during the process acquires a profile consistent with that of the stationary solution of the FDM+BH eigenvalue problem. These results indicate that BHs can indeed act as focal points for FDM core condensation. Furthermore, we find that the central density of the resulting FDM core depends on the mass of the BH, which due to its permanent motion relative to the FDM core during the evolution, produces a smaller core density for bigger BH masses; in this way the BH mass is a parameter leading to a new diversity of central FDM core densities. As a collateral result, for our analysis we revised the construction of stationary solutions of FDM+BH and found a phenomenological formula for the FDM density that can be used to fit FDM cores around BHs.
@@ I Introduction
Fuzzy Dark Matter (FDM) assumes dark matter is an ultralight boson with mass of order $10^{-23}-10^{-21}$eV, whose behavior differs from that of cold dark matter at galactic scales, in particular it promotes core formation and shows wave-like phenomena that produce a particular distribution of matter at galactic core and halo scales Chavanis (2015); Niemeyer (2020); Hui (2021); Ferreira (2020). The structures formed by this particle are cores surrounded by envelopes with profiles similar to those of the CDM, the core being an essential fingerprint of the model determined from structure formation simulations (e.g. Schive et al. (2014a); Mocz et al. (2017); Veltmaat et al. (2018); Mocz et al. (2019); May and Springel (2021); Schwabe and Niemeyer (2022)).
@@
Core formation has been studied in more detail at small scale simulations in basically two settings, one is the multi-merger of cores, for example in Schive et al. (2014); Mocz et al. (2017); Schwabe et al. (2016); Álvarez-Rios et al. (2023); Álvarez-Rios and Guzmán (2024) that leads to a simple construction of core-halo structures. Another approach uses the kinetic relaxation as explanation and simulation of core formation out of random initial conditions Levkov et al. (2018); Chen et al. (2021). These two methods help studying the core condensation process of bosonic configurations, their mass-growth and the envelope profile of their halos using small domain simulations Eggemeier and Niemeyer (2019); Chen et al. (2023); Purohit et al. (2023); Chen and Zhang (2024).
@@
An essential ingredient introduced in FDM phenomenology is the presence of black holes and their behavior within FDM cores. For example, in Chavanis (2019, 2020) the FDM core density profile properties are studied under various regimes of the boson gas and scenarios that include a black hole, in Wang and Easther (2022) the interaction of FDM and the BH is studied, in particular the dynamical friction and the drag of the wake behind the hole while moving, in Boey et al. (2024) also the dynamical friction of FDM cores in BH is studied in merger scenarios, in Koo et al. (2024) also the binary black hole merger moving inside FDM cores is studied in the context of the final parsec problem of the merger. Other studies reveal the impact that FDM cores have at boosting supermassive black hole growth Chiu et al. (2025). In gravitational wave related contexts, the binary black hole within FDM is studied to estimate the extraction of angular momentum due to dynamical effects of clumps of dark matter within the core Bromley et al. (2024), and more recently the scattering of FDM by the binary is analyzed Tomaselli (2025). In El-Zant et al. (2020) the authors study the ejection of SMBHs due to the superposition of modes and the accumulation of random walk effects within FDM halos. Also in Lancaster et al. (2020) the motion of a massive point particle in a FDM environment that includes its granules is studied, while in Boey et al. (2024) the collision between an FDM core and a SMBH is analyzed. These studies consider the black hole to be Newtonian, and focus mainly in dynamic effects of moving black holes. In a relativistic context, the coexistence and phenomenology of black hole with ultralight bosonic dark matter has also been analyzed, for example in Cardoso et al. (2022a); Traykova et al. (2023); Ravanal et al. (2023); Boudon et al. (2023); Traykova et al. (2023) various effects of black hole dynamics on the scalar field are studied, including dynamical friction, while in Bamber et al. (2023); Aurrekoetxea et al. (2024a) the analysis centers on the potential detection of ultralight dark matter via effects on gravitational wave phenomenology, including the case with self-interaction Aurrekoetxea et al. (2024b). In cosmological scenarios, studies include for example the nucleation of scalar clouds Hertzberg et al. (2020).
@@
In this work we study the effects of a BH during the FDM core-condensation, for which we follow the kinetic relaxation method developed in Levkov et al. (2018); Chen et al. (2021), that uses random initial conditions, so that the granularity of the FDM distribution develops in the presence of the BH since initial time. In our analysis we use the Newtonian version of a spherically symmetric BH and add its gravitational effects to the Scrödinger-Poisson system of equations that rules the dynamics of the FDM+BH system, while we ignore the partial accretion of wave dark matter, which depends on the wave-length and thickness of wave-packets approaching the black hole, as demonstrated with full non-linear numerical relativity in Guzmán and Lora-Clavijo (2012); Guzmán et al. (2021), as well as other non-linear accretion effects in non-symmetric scenarios Cardoso et al. (2022b).
We find that the FDM distributes around the BH, where it starts the process of condensation and that its density approaches that of stationary solutions of the FDM+BH eigenvalue problem. In order to show this, we revise the construction of stationary solutions of this system following the ideas in Davies and Mocz (2020) and build a phenomenological formula usable to fit FDM distributions during the evolution of the system. We also use these solutions to test our evolution code.
During the condensation process we notice that the interaction of the FDM structure with the BH, that back reacts to the dynamics of the FDM, produces an oscillatory motion of the BH as described in Boey et al. (2024), in this case in a granular non-smooth FDM distribution. We also find that the relative motion of the BH with respect to the core influences the final FDM distribution and flattens the central core density. We consider this effect is due to the BH scattering of FDM while moving within the core; in order to support this idea, we simulated the case in which the BH does not back-react to the gravity sourced by the FDM and observe that in such case there is no central density decrease.
The paper is organized as follows. In Section II we describe the set of equations used to simulate the system, along with a description and tests of our code. In Section III we describe the simulations carried out and their diagnostics. Finally, in Section IV we draw some conclusions of our analysis. We include three appendices, one in which we revise the stationary solution of the FDM+BH eigenvalue problem and construct a phenomenological formula that mimics the FDM distribution; a second one in which we evolve the system without BH, that serves as a test of our code and for comparison with our findings; a third one in which we study the test-field regime that serves to understand better the role played by the BH motion.
II Equations and numerical methods
II.1 Evolution Equations
@@
The system of equations that rule the dynamics of the FDM+BH system is the following Schrödinger-Poisson (SP) system of equations:
@@
where $m$ is the boson mass, $\rho=m|\Psi|^{2}$ is the bosonic gas density, $\bar{\rho}$ its mean density, $V$ is the gravitational potential due to the FDM and the BH, while $V_{FDM}$ is the potential only due to the FDM. We approximate the BH with a Gaussian density distribution:
$\rho_{BH}=M_{MH}\delta(\vec{x}-\vec{x}_{BH})\simeq CM_{MH}e^{-|\vec{x}-\vec{x}_{BH}|^{2}/2\epsilon^{2}}$ (5)
where $C$ is such that the integral of the distribution is the black hole mass $M_{BH}$. The total density of the system is $\rho_{T}=\rho+\rho_{BH}$ so that the BH contributes to the total gravitational potential of the system, and finally, $\vec{x}_{BH}$ is the position of the black hole.
@@
In order to solve the SP system above, we use the transformations $t=t_{0}\tilde{t},\vec{x}=x_{0}\tilde{\vec{x}},V=V_{0}\tilde{V},\Psi=\Psi_{0}\tilde{\Psi}$ and $\rho=\rho_{0}\tilde{\rho}$, where $t_{0}=\frac{x_{0}^{2}m}{\hbar}$, $V_{0}=\left(\frac{\hbar}{mx_{0}}\right)$, $\Psi_{0}=\frac{\hbar}{\sqrt{4\pi Gm^{3}}x_{0}^{2}}$, $\rho_{0}=\frac{\hbar^{2}}{4\pi Gm^{2}x_{0}^{4}}$, that leave the SP system in dimensionless code units that only depend on the length scale parameter $x_{0}$. We solve the dimensionless problem using the code CAFE-FDM Álvarez-Rios and Guzmán (2022), that implements a pseudo-spectral method to discretize spatial derivatives, an RK4 scheme for the evolution of the wave function, and the two Poisson equations are solved using the FFT method. All simulations are carried out in a periodic cubic domain of side $L$ in code units, with resolution $\Delta x$. For the BH distribution we use $\epsilon=0.1\Delta x$ because it gives the same results as in Boey et al. (2024), where a BH is represented thorough its gravitational potential $V_{\bullet}=-\frac{GM_{BH}}{\sqrt{max(r^{2},\epsilon^{2})}}$ and used within Schrödinger equation to simulate scenarios with FDM and BH in isolated domains; in fact we define the BH in terms of the distribution $\rho_{BH}$ because it works in periodic domains as well.
II.2 Initial conditions
@@
The kinetic relaxation leading to core condensation uses random initial conditions for the bosonic gas, that eventually lead to the collapse of overdensities that in turn promote the bosonic gas condensation. The systematic study of this collapse process is understood from local simulations defined in Levkov et al. (2018); Chen et al. (2021), where various distributions in the momentum domain are proposed. In our analysis we use a Gaussian distribution $\Psi(\vec{p})=Ae^{-p^{2}/2}e^{iS}$ in the momentum space, with $S$ a random phase in the range $[0,2\pi]$ at each point of the momentum space, where $A$ is a normalization factor. In order to validate our methods and code, we run one of the simulation in Chen et al. (2021), specifically on a box of side $L=18$, with resolution $\Delta x=L/128$ and total FDM mass in the numerical domain $M_{FDM}=M=1005.3$ in code units as in Chen et al. (2021). As an example in physical units, we use the scenario where cores of FDM are shown to boost the growth of supermassive black holes Chiu et al. (2025), there the boson mass is $m=10^{-22}$eV, and the average FDM density is $\bar{\rho}\sim 1.477\times 10^{8}$${}_{\odot}/$kpc3. The integral in our domain gives $M=2.29\times 10^{10}$, whereas the numerical resolution and domain size for such boson mass are $\Delta x=42$pc and $L=5.3749$kpc. Under these conditions, the smallest black hole mass in our analysis below will be $M_{BH}=M/256=8.96\times 10^{7}$ and the most massive one $M_{BH}=M/32=7.168\times 10^{8}$. Another important part of information is the momentum space distribution, notice that the total mass of FDM in the domain defines the value of the normalization constant $A$ of the momentum space distribution. Moreover, we use a distribution of width $\sigma=1$, at one sigma radius $p=1$, the associated velocity is $\sim 65$km/s, which is cold.
In what follows we use code units and use the numerical setting above as default in our simulations involving BHs. A final specification of initial conditions is that for the BH the initial condition consists in setting its location at the center of the domain.
II.3 Tests.
As a first test, we reproduce the core condensation of FDM with kinetic relaxation using the initial conditions described above, that leads to results consistent with those in Chen et al. (2021). The test in contained in Appendix A and also serves to compare the condensation process with the cases with a Black Hole.
As a second test of our implementation, we evolve models of the stationary solutions of the FDM+BH eigenvalue problem developed in Davies and Mocz (2020) and revised in Appendix B. We show that the solutions are stable and that our phenomenological formula is usable.
III Simulations and analysis
III.1 Simulations
@@
We performed 32 simulations with the same type but different seed to generate random initial conditions for the FDM and various black hole masses $M_{BH}=M/256,M/128,M/64,M/32$, which lie within the range of masses used in studies of dynamical friction Wang and Easther (2022) and those of PBHs used for nucleation of axion stars in Hertzberg et al. (2020). Different initial seeds for initial conditions lead to different evolutions of the FDM in the presence of a BH, however the results of the simulations can be classified into two sets: a) a minicluster of FDM forms near the BH and its further collapse, relaxation and core condensation happen around the BH all the way, b) a minicluster of FDM forms beyond the tidal radius of the BH and its further evolution includes the merger with the BH where the relaxation and condensation into a core happens. Among these two scenarios the second one is the non-trivial case that better shows how the BH acts as a condensation point for FDM density. In order to illustrate the evolution of this generic case, we use a particular seed to generate the random phase of FDM in Fourier space that in turn generates the initial conditions of the FDM, and study its evolution using the four black hole masses. The evolution of these four simulations is presented in Figure 1, that contains snapshots of the FDM density and the Black Hole position. At early times by $t\sim 7$ a minicluster forms near the bottom-right corner, later on this cluster merges with the black hole, following different trajectories for each black hole mass, and finally the condensation takes place at the black hole location. Animations of these simulations appear as Supplementary Material Palomares-Cávez et al. .
@@
Figure 1: Snapshots of the density projected on a plane that contains the position of the black hole and is parallel to the $xy-$plane of the numerical domain. The black hole position is represented with a yellow cross at different times for simulations with $M_{BH}=M/256,M/128,M/64$ and $M/32$. The evolution illustrates the formation of a minicluster near the bottom right corner. Later on the minicluster and the BH merge and start traveling together while the BH position oscillates with respect to the core, which can be seen in the movies in the supplementary material. By $t\sim 60$ the FDM approaches condensation as seen in the analysis below.
III.2 Condensation.
@@
In order to see if condensation happens, we monitor the evolution of $\rho_{max}$ and the results are shown in Figure 2, where we notice that the condensation takes place, the density grows with a power law of $t$ starting after $t\sim 30$, which indicates the beginning of condensation, however later on the maximum density decreases or stabilizes instead of growing up as in the pure FDM scenario in Levkov et al. (2018); Chen et al. (2021). The decrease of $\rho_{max}$ is smaller for smaller black hole masses and can be understood in terms of the relative motion between the BH and the core. That is, the black hole, by being in motion, drags the FDM around it and prevents the density from increasing, with the effect being stronger for larger black hole masses. Thus, the presence of the black hole has the effect of flattening the FDM core density in the long term.
We contrast this density behavior with that it the test field case developed in Appendix C, in which the Black Hole does not back-react to the dynamics of the FDM and remains at a fixed position. In this scenario the FDM is attracted toward the hole and condensates around it, the black hole does not move, there is no re-heating due to the FDM dynamics and the BH does not scatter the FDM around. The comparison between Figures 2 and 12 explains this effect.
@@
Figure 2: Evolution of $\rho_{max}$ for the various black hole masses $M_{BH}=M/256,M/128,M/64,M/32$. Notice that for the less massive BHs the density continues growing until $t\sim 60$ where it stabilizes. A difference between small BH masses and the case $M_{BH}=0$ shown in Fig 7, is that $\rho_{max}$ stabilizes with a BH, whereas it keeps growing when there is not BH in this time window, although as we will see below, the central density for the cases $M_{BH}=0,M/256$ and $M/128$ are very similar. For the most massive BHs the maximum density decreases, which is an indication of the effects of the presence of the BH. For comparison of the evolution without BH see the simulations in Appendix A.
III.3 Relative motion.
@@
In Figure 3 we show the distance from the point of maximum density to the Black Hole, for the four values of $M_{BH}$. The non-linear behavior and the dynamics of the granularity lead to a behavior that does not have a trend, since the motion of the black hole can be reheated differently since the different four mergers occur under different conditions, at different places and with different merging velocities, as can be seen in the animations and in Figure 1.
@@
Another part of the analysis is the velocity of the BH. In Figure 4 we show the $x-$component of the black hole’s velocity, which shows the oscillatory behavior studied under smoother FDM distributions in Boey et al. (2024). Moreover, for the cases $M_{BH}=M/128$ and $M/64$ the reheating pulses can be observed in the velocity as also described in Boey et al. (2024). Unlike in Boey et al. (2024) the oscillatiory motion is less regular because we do not assume an initially smooth spherically symmetric core, but deal with the core resulting from the collapse of FDM, with all its granular multipolar components and time dependence, since the formation time.
@@
Figure 4: The $x-$component of the BH velocity for $M_{BH}=M/256/,M/128,M/64,M/32$. The oscillations are consistent with those proposed in Boey et al. (2024) for mergers of smooth cores and BHs. In our case the core has all its granules and time dependency since the beginning. Notice that for $M_{BH}=M/128,M/64$ the modulation of the velocity amplitude is consistent with the reheating of the BH motion. The behavior of the $y$ and $z-$components of the velocity is similar. The animations in the Supplementary Material illustrate the whole picture of the process.
III.4 Core fitting.
@@
Even though there is dynamics, we fit the resulting density profile of the FDM, not only at a fixed time but averaged over a time window using a phenomenological model for the density, that mimics the solution of the FDM-BH stationary eigen-problem in Davies and Mocz (2020), as described in Appendix B; the FDM density fitting uses formulas (11)-(13). The result shows the possibility to fit the configuration with an equilibrium profile of the FDM-BH eigen-problem. The fittings are presented in Figure 5, that we obtain from averaging the density profile in the time window $t\in[70,100]$, which is a lapse when, according to Fig. 2, the maximum density profile ends growing during condensation for the smallest black holes. Notice that the central core density is smaller for bigger BH mass, whereas the central core density converges to the case without BH when $M_{BH}$ becomes smaller, in fact the cases $M_{BH}=M/256$ and $M/128$ have a profile similar already to that of the case $M_{BH}=0$, which is an interesting consistency check of the simulations. This does not mean that small BHs play no role at all, notice that for all BH masses the minicluster truly moves and merges with the BH; this shows that even if small, BHs still play the role of condensation points.
@@
Figure 5: Density fitting of the angularly and time averaged density of FDM around the BH, using the phenomenological formula (11), for the simulations with $M_{BH}=M/256,M/128,M/64,M/32$. The time interval used to average density profiles is $t\in[70,100]$. Notice that the more massive the black holes, the smaller the central core density, which is consistent with the notion of a Black Hole moving within the core while dragging FDM around and preventing the density from growing, which is consistent with the results in Figure 2. We enclose the case for $M_{BH}=0$ to show that when the mass of the BH approaches zero, the core approaches the case without BH.
IV Conclusions